OpenStudy (anonymous):

In a farm, there are 600 banana trees. The number of tress in each row is 10 less than twice the number of rows. How many trees are in each rows?

OpenStudy (lgbasallote):

let x = number of rows that means 600/x = number of trees per row it says the number of trees per row is also 10 less than twice the number of rows so 600/x = 2x - 10 make sense? now multiply all terms by x 600 = 2x^2 - 10x turn it into general form of quadratic equation 2x^2 - 10x - 600 = 0 divide everything by 2 (to make it simpler) x^2 - 5x - 300 = 0 factor it out (x-20)(x+15) = 0 solve for x x = 20; x = -15 obviously it's impossible to be negative so take the positive root does that help?